Rank these engines on the basis of their designed
Answer #1

Output power

\(P=f\left(Q_{\text {in }}-Q_{\text {out }}\right)\)

\(P_{1}=200 \mathrm{~Hz} \times(1500 \mathrm{~J}-500 \mathrm{~J})=200000 \mathrm{~W}\)

\(P_{2}=150 \mathrm{~Hz} \times(3000 \mathrm{~J}-1500 \mathrm{~J})=225000 \mathrm{~W}\)

\(P_{3}=200 \mathrm{~Hz} \times(1000 \mathrm{~J}-200 \mathrm{~J})=160000 \mathrm{~W}\)

\(P_{4}=50 \mathrm{~Hz} \times(2000 \mathrm{~J}-500 \mathrm{~J})=75000 \mathrm{~W}\)

\(P_{5}=50 \mathrm{~Hz} \times(4000 \mathrm{~J}-3000 \mathrm{~J})=50000 \mathrm{~W}\)

\(P_{6}=100 \mathrm{~Hz} \times(2000 \mathrm{~J}-1000 \mathrm{~J})=100000 \mathrm{~W}\)

Hence

\(P_{2}>P_{1}>P_{3}>P_{6}>P_{4}>P_{5}\)

Engine order from largest power output to smallest power output.

\(2,1,3,6,4,5\)